Please help me answer this problem:
"The difference in the length of two segments is 15 and the ratio is 3:8. What are the lengths of the segments?"
Thank You.

Answer by thejackal(72)

let one segment be n, therefore the other segment is n + 15
the ratio is given thus 3/8 = n/n+15;
now cross multiply to get
3(n+15) = 8(n)
simplify
3n + 45 = 8n
45 = 5n
9 = n
therefore the longer segment is 9 + 15 = 24
QED